If the parabola y-2=a(x-30^2 goes through the point (2,0), what is the value of a?
Answer by lwsshak3(11628)

If the parabola y-2=a(x-30^2 goes through the point (2,0), what is the value of a?
..
I will assume you meant to write given equation as follows:
y-2=a(x-30)^2
plug in (x,y) coordinates from given point (2,0) into equation
0-2=a(2-30)^2=a(28)^2=784a
a=-2/784=-.0026
equation: y=-.0026(x-30)^2+2
This is a parabola which opens downward with vertex at (30,2)
a=-.0026